Present and future value tables of $1 at 3% are presented below:

N FV $1 PV $1 FVA $1 PVA $1 FVAD $1 PVAD $1
1 1.03000 0.97087 1.0000 0.97087 1.0300 1.00000
2 1.06090 0.94260 2.0300 1.91347 2.0909 1.97087
3 1.09273 0.91514 3.0909 2.82861 3.1836 2.91347
4 1.12551 0.88849 4.1836 3.71710 4.3091 3.82861
5 1.15927 0.86261 5.3091 4.57971 5.4684 4.71710
6 1.19405 0.83748 6.4684 5.41719 6.6625 5.57971
7 1.22987 0.81309 7.6625 6.23028 7.8923 6.41719
8 1.26677 0.78941 8.8923 7.01969 9.1591 7.23028
9 1.30477 0.76642 10.1591 7.78611 10.4639 8.01969
10 1.34392 0.74409 11.4639 8.53020 11.8078 8.78611
11 1.38423 0.72242 12.8078 9.25262 13.1920 9.53020
12 1.42576 0.70138 14.1920 9.95400 14.6178 10.25262
13 1.46853 0.68095 15.6178 10.63496 16.0863 10.95400
14 1.51259 0.66112 17.0863 11.29607 17.5989 11.63496
15 1.55797 0.64186 18.5989 11.93794 19.1569 12.29607
16 1.60471 0.62317 20.1569 12.56110 20.7616 12.93794

  
Today, Thomas deposited $120,000 in a 2-year, 12% CD that compounds quarterly. What is the maturity value of the CD?

Multiple Choice

  • $127,200.

  • $253,354.

  • $232,028.

  • $152,012.
Answer #1

Maturity value of the CD

= 120,000 * FV of $1 at N=4

= 120,000 * 1.26677

= 152,012

Option D is the answer